Compute the amount of the trade discount and the net price using the discount method.

List

Trade

Discount

Net

Price

Discount

Amount

Price

$2,672

35%

__________

__________

Group of answer choices

Discount Amount: $935.20; Net Price: $1,736.80

Discount Amount: $295.20; Net Price: $1,344.80

Discount Amount: $158.40; Net Price: $561.60

none of these

--------------------------------------------------------------------

List

Trade

Discount

Net

Price

Discount

Amount

Price

$1,640

18%

__________

__________

Compute the amount of the trade discount and the net price using the discount method

Group of answer choices

Discount Amount: $295.20; Net Price: $1,344.80

Discount Amount: $935.20; Net Price: $1,736.80

Discount Amount: $158.40; Net Price: $561.60

none of these
Answer #1

1. Discount amount = 2672*35% = 935.20

Net price = 2672 - 935.20 = 1736.80

Option A

2. Discount amount = 1640*18% =295.20

Net price = 1640 - 295.20 = 1344.80

Option A